OpenStudy (anonymous):

How do you write this in the quadratic equations in vertex form? y=x^2-4x+6

OpenStudy (anonymous):

You need to use the method called "Completing the square". Given \[y=x^2 +bx +c\] you can complete the square by \[y=x^2 +bx +\left(\frac{b}{2}\right)^2-\left(\frac{b}{2}\right)^2 + c\] note that by adding and subtracting the (b/2)^2 term I haven't changed the equation at all (like adding 5 and subtracting 5 doesn't change a value) But we know that \[\left(x+\frac{b}{2}\right)^2= x^2 + \frac{b}{2}x+\frac{b}{2}x +\left(\frac{b}{2}\right)^2 \] \[\left(x+\frac{b}{2}\right)^2= x^2 + 2\frac{b}{2}x+\left(\frac{b}{2}\right)^2 \] \[\left(x+\frac{b}{2}\right)^2= x^2 + bx+\left(\frac{b}{2}\right)^2 \] so \[y=x^2 +bx +\left(\frac{b}{2}\right)^2-\left(\frac{b}{2}\right)^2 + c\] therefore becomes \[y=\left(x+\frac{b}{2}\right)^2-\left(\frac{b}{2}\right)^2 + c\] which is essentially in the vertex form of y=k(x-p)^2+q where k=1, p=-(b/2), and q= c-(b/2)^2 Note that if you start with something like \[y=ax^2 + 4x+13\] that you should divide out by a first so you have \[\frac{y}{a}= x^2 + \frac{4}{a}x+\frac{13}{a}\] do the process and then at the END multiply through by a. That will give you the \[y=k(x-p)^2+q\] which is the vertex form of a quadratic that gives the vertex at (p,q) the axis of symmetry at x=p and whether it is open up (k is positive) or open down (k is negative)

OpenStudy (anonymous):

So for example if I had \[y=x^2 -100x +1\] I would then write \[y=x^2 -100x +\left(\frac{-100}{2}\right)^2-\left(\frac{-100}{2}\right)^2 +1\] \[y=x^2 -100x +\left(-50\right)^2-\left(50\right)^2 +1\] and then \[y=(x-50)^2 -2500 +1\] \[y=(x-50)^2 -2499\] So the vertex would be at (50,-2499), an axis of symmetry at x=50, and it would be open up.